Subject: 3.5-m Users Committee minutes 12/10/07

From: Bruce Gillespie

Submitted: Wed, 9 Jan 2008 10:05:44 -0700

Message number: 1080 (previous: 1079, next: 1081 up: Index)

                    APO 3.5-m Users Committee Phonecon, 12/10/07

Attending:  Suzanne Hawley, Bruce Gillespie, Scott Anderson, Ed Turner  
(for Michael Strauss), Remy Indebetouw, John Bally, Al Harper, Jon  
Holtzman, Russet McMillan

Absent: Jon Fulbright, Mark Klaene, Michael Strauss

Minutes taken by Bruce Gillespie


User feedback, comments from institutional representatives:

Princeton (Turner) - Ed said that he didn't have any user feedback to  

UVa (Indebetouw) - Remy said the Virginia users are happy with their  
recent observing runs.

UW (Anderson) - Scott had nothing new to report from the user side.

CU (Bally) - John had no new user feedback to report.

JHU (Fulbright) - No report.

NMSU (Holtzman) - Jon had no user feedback about recent observing.

UChicago (Harper) - Al reported that Chicago users had recently had  
good observing runs.


Discussion of telescope/instruments report - Gillespie:

The report is given below.  Bruce walked through the highlights, to  
which there was no substantive discussion:


3.5-m Telescope, Instruments, and CIF Projects Highlights, 11/15/07  
through 12/05/07

                           Mark Klaene & Bruce Gillespie

0) Overview

Sporadic winter weather has finally started to move in, mostly in the  
form of occasional rain, near-freezing temperatures, and wind.   
Observing-time losses due to weather increased slightly over the  
previous month, but overall the skies have been generally usable for  
observing.  The site is preparing for the holiday break when much of  
the day staff will be on leave or standby status.  SDSS-III  
construction continued at a reasonable pace without interference to  
nighttime observing.

1) Telescope

We are optimistic that recent problems with tertiary rotation may have  
finally been solved.  After extensive diagnostic effort, it was  
discovered that a scale bit in the tertiary rotation controller was  
not being correctly reset by an electrical jumper.  Since this problem  
was found and corrected, there has been no further misbehavior of the  
tertiary rotation system.

2) Instruments

DIS, SPIcam, and the echelle have all operated smoothly throughout the  
period, with no reported problems.  The DIS spare Cryotiger compressor  
will be repaired and returned to the site by mid-January.  The  
performance of the NIC-FPS ion pump has returned to nominal levels  
during the past month; the ion pump is being monitored with no near- 
term action currently planned.  Also for NIC-FPS, CU scientists  
suspect they may know the cause of the increased detector noise that  
manifested since the upgrade to 16-channel read capability, and are  
fabricating a new fan-out board that would be installed next summer.   
NIC-FPS etalon testing continues towards reaching the requisite plate  
parallelism.  The new focal reducer for Agile has been AR-coated and  
reinstalled in the instrument.  Agile characterization tests were  
weathered out and a retry is planned for the next available  
opportunity.  The visiting instruments (CorMASS, GFP, and APOLLO) are  
all operational.

3) CIF projects

The designs for Agile field rotation and its move to TR2 port are  
nearly complete.  ALso, good progress was made on the new axis  
controller for the instrument rotation axis--the system is currently  
undergoing operational testing at night.


NIC-FPS status - Bally:

John talked about the problems with the NIC-FPS etalon, namely  
difficulties in obtaining plate parallelism.  Corey Wood will be  
coming to APO to test the etalon with an artificial star.  One  
possible cause of the problems may be related to the length of cable  
runs to the instrument.  Also, a new, better shielded, fan-out board  
for the detector is being fabricated and will be installed the next  
time the instrument dewar is opened, with the hope of reducing read  
noise.  The coronagraphic mode is now working, and uses new masks.   
There is a general problem with residual images when bright objects  
are observed; the residual images take ~hours to clear from the  
detector.  It may be advisable for programs that have bright target  
objects to take these observations at the end of the night, although  
this would add a burdensome constraint to telescope scheduling.  Al  
Harper suggested that experiments could be undertaken to mitigate the  
residual image problem, and that other groups that are using the  
Rockwell RG series detectors (e.g., the JWST group at GSFC) may have  
already found workarounds.  John also said that a new "random" dither  
pattern (+/- 4 or 5 arcseconds) is highly recommended.  It is unclear  
if this is currently supported within TUI.

ACTION:  User Committee members should find out if their NIC-FPS users  
are having problems with residual images caused by bright targets.


Time trades with WIYN - Hawley:

Suzanne reminded us that she has been working with George Jacoby to  
arrange for time-swapping between WIYN and our 3.5-meter telescope.   
WIYN may have its One-Degree Imager (ODI) coming on line in the next  
year or two which would be a unique capability that we don't offer.   
The WIYN users appear to be most interested in using DIS.  Anyone  
interested in access to WIYN should get in touch with Suzanne.  The  
time trades would begin in about six to eight months.


User documentation - Hawley/Holtzman:

Jon said that he is still taking user feedback on the new  
instrumentation web pages.  The link to the new pages will be the  
default in the next month or so; the old pages will be archived and  
their links kept.



Ed Turner thanked the users who provided input to him regarding the  
reappointment of Suzanne as Director of the 3.5-m telescope.  The ARC  
Board of Governors was pleased that Suzanne accepted the reappointment  

Suzanne mentioned that an NMSU grad student has begun Observing  
Specialist training.  The idea is that this person would be available  
for emergency fill-in situations on short notice.


[open from previous months]:

ACTION:  Committee members to look at the new user webpages, and  
report back to Jon Holtzman and Gabrelle Saurage suggestions for  
content and organization.

   STATUS:  Open, Jon and Gabrelle waiting for more feedback.

ACTION:  John Bally will write a section with figures on Fowler  
sampling for inclusion in the NIC-FPS user documentation.

   STATUS:  Open.

[new actions from this month]:

ACTION:  User Committee members should find out if their NIC-FPS users  
are having problems with residual images caused by bright targets.


Next meeting:

The next Users Committee phonecon will be on Monday, 14 January, at  
8:30 AM Pacific Time.  The agenda and other materials will be sent to  
the committee members during the preceding week.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 1080 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to